### v2

@@ 7.2.2 Outcomes without an infinitely occurring vertex
We now study NE outcomes in which no vertices occur infinitely often. These outcomes are specific to infinite arenas. We first present how to derive well-structured NE outcomes from general NE outcomes. We obtain NE outcomes with an aperiodic infinite decomposition. We then present a variant of the previously used two-phase mechanism for these outcomes. While we leave the first phase is unchanged, finite memory no longer suffices to implement a decomposition-based approach for the second phase. Instead, we use a looser (but sufficient) variant of decomposition-based strategies to construct NEs.
@@
We first provide a counterpart to Lemma 7.4 for this case. Let $\pi^{\prime}\in\mathsf{Plays}(\mathcal{A})$ be the outcome of an NE from $v_{0}\in V$ such that no vertex occurs infinitely often in $\pi$. We observe that $\pi^{\prime}$ need not be a simple play: the objectives satisfied by $\pi$ may require visiting targets that are not connected by a simple play or a lasso. Therefore, simple plays do not suffice for our purpose, and we need memory to construct the outcome (even without considering the stability of the NE).
We construct well-structured NE outcomes that can be obtained with two memory states. More precisely, we derive an NE outcome $\pi$ from $\pi^{\prime}$ that has an infinite simple segment decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ such that $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, for all odd (resp. even) $j\neq j^{\prime}$, no vertex of $\mathsf{sg}_{j}$ occurs in $\mathsf{sg}_{j^{\prime}}$ and, for all $j\geq 1$, no targets of players whose objective is not satisfied by $\pi^{\prime}$ occurs in $\mathsf{sg}_{j}$. The condition on segments with the same parity implies that $\pi$ is the outcome of a strategy profile with memory $\{1,2\}$: intuitively, memory state $1$ is used when following an odd segment and $2$ is used when following an even segment.
@@
We define a finite-memory NE from $v_{0}$ whose outcome is $\pi$ via the two-phase mechanism previously used in the proof of Theorem 7.6. The first phase is defined identically to the previous section (see the proof of Theorem 7.6). For the second phase, we use a variation of the partially-defined Mealy machines of Section 5.3. In the definitions of Section 5.3, for each segment index $j$ and each memory state of the form $(\mathcal{P}_{i},j)$ , we refer to the simple segment $\mathsf{sg}_{j}$ to define memory updates and next moves. This approach no longer yields a finite-memory strategy with the infinite decomposition $\mathcal{D}$. Instead of having a group of memory states for each segment, we use a group of memory states for each parity. More precisely, we use memory states of the form $(\mathcal{P}_{i},1)$ for all odd segments and memory states of the form $(\mathcal{P}_{i},2)$ for all even segments besides $\mathsf{sg}_{0}$. When the end of an odd (resp. even) segment is reached in a memory state of the form $(\mathcal{P}_{i},1)$ (resp. $(\mathcal{P}_{i},2)$), the memory is updated to a state of the form $(\mathcal{P}_{i^{\prime}},2)$ (resp. $(\mathcal{P}_{i},1)$). We obtain a well-defined next-move function because segments of $\mathcal{D}$ with the same parity traverse pairwise disjoint sets of vertices.
@@
The update function $\mathsf{up}$ is defined as follows. For all $\ell<r$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$. We postpone the definition of $\mathsf{up}(v_{r},v_{r})$. For all $\ell\leq r$ and $v\neq v_{\ell}$, if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$ (i.e., $v\in V_{I}$), we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$, and otherwise the update is arbitrary. Let $p\in\llbracket{}2\rrbracket{}$. For all states of the form $(\mathcal{P}_{i},p)\in M$ and all $v\in S_{p}$, we let $\mathsf{up}((\mathcal{P}_{i},p),v)=(\mathcal{P}_{i^{\prime}},p^{\prime})$ where (a) $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$, and (b) $p^{\prime}=p$ if $v\notin L_{p}$, and otherwise we set $p^{\prime}=3-p$ (i.e., if $p=1$, it becomes $2$ and vice-versa). For all states of the form $(\mathcal{P}_{i},p)\in M$ and all $v\in V\setminus S_{p}$, we let $\mathsf{up}((\mathcal{P}_{i},p),v)=\mathcal{P}_{i}$. We let $\mathsf{up}(v_{r},v_{r})=\mathsf{up}((\mathcal{P}_{i},1),v_{r})$ where $i\in I$ is arbitrary (the definition does not depend on $i$ because $v_{r}\in S_{1}$). Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$. We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq r$. If $v=v_{\ell}$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$ if $\ell\neq r$ and let $\mathsf{nxt}_{i}(v_{r},v_{r})$ be the second vertex of $\mathsf{sg}_{1}$ (which exists because $\mathsf{sg}_{1}$ is non-trivial). Assume that $v\neq v_{\ell}$. If $\ell\geq 1$ and $v_{\ell-1}\in V_{I}$, we let $i^{\prime}\in I$ such that $v\in V_{i^{\prime}}$ and set $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$. We let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases.
@@
We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},p)$. Fix $i^{\prime}\in I$ and $p\in\llbracket{}2\rrbracket{}$. Assume first that $v\in S_{p}\setminus L_{p}$. There is a unique $j\in 2\mathbb{N}+p$ such that $v$ occurs in $\mathsf{sg}_{j}$ (it is unique because all segments with the same parity traverse disjoint sets of vertices). We set $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ to the vertex following $v$ in $\mathsf{sg}_{j}$. Next, assume that $v\in L_{p}$. There is a unique $j\in 2\mathbb{N}+(3-p)$ such that $v=\mathsf{first}(\mathsf{sg}_{j})$. We let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ be the second vertex of $\mathsf{sg}_{j}$ (which exists because $\mathsf{sg}_{j}$ is non-trivial). If $v\notin S_{p}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ be arbitrary if $i^{\prime}=i$. Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)$ is arbitrary otherwise.
@@ 7.3.1 Well-structured Nash equilibrium outcomes
In co-Büchi games, these deviations can be harmful: if we use a fully decomposition-based approach to defining strategies from NE outcomes, plays resulting from profitable in-segment deviations may violate some objectives satisfied by the original NE outcome (Example 7.1). This motivates the use of a two-phase approach in co-Büchi games: we adopt a decomposition-based behaviour once in-segment deviations can no longer be harmful to winning players. Formally, we show that from any NE outcome $\pi^{\prime}$ in $\mathcal{G}$, we can derive an NE outcome $\pi$ from $\mathsf{first}(\pi^{\prime})$ that is a simple play or a simple lasso such that $\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi)$, and there are no plays $\pi^{\prime\prime}$ from $\mathsf{first}(\pi{r^{\star}})$ such that $\mathsf{SatPl}_{\mathcal{G}}(\pi)\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$ and $\pi^{\prime\prime}$ uses only vertices appearing in $\pi{r^{\star}}$ where $r^{\star}$ is the smallest $r\in\mathbb{N}$ such that no targets of players in $\mathsf{SatPl}_{\mathcal{G}}(\pi)$ appear in $\pi{r}$. The prefix $\pi{r^{\star}}$ corresponds to the first phase and the suffix $\pi{r^{\star}}$ to the second phase of our approach, and the condition imposed on $\pi$ guarantees the absence of profitable deviations. To construct $\pi$, we proceed by induction: we repeatedly simplify $\pi^{\prime}$ into well-chosen simple lasso or simple plays until we obtain a suitable play. The following proof provides a formal construction.
@@
Assume that $\rho^{(j)}=v_{0}v_{1}\ldots$ is well-defined and let $r_{j}=\min\{r\in\mathbb{N}\mid\forall\ell\geq r,\,\forall i\in\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j)}),\,v_{\ell}\notin T_{i}\}$. We define $\pi^{(j)}$ as a combination of a history $\mathsf{sg}_{j}$ and a play $\mathsf{sg}_{j}^{\prime}$. We distinguish two cases for the suffixes. If $\rho^{(j)}{r_{j}}$ is a simple play, we let $\mathsf{sg}_{j}^{\prime}=\rho^{(j)}{r_{j}}$. Otherwise, we let $\mathsf{sg}_{j}^{\prime}$ be a simple lasso from $\mathsf{first}(\rho^{(j)}{r_{j}})$ such that all states that occur in $\mathsf{sg}_{j}^{\prime}$ occur in $\rho^{(j)}{r_{j}}$. For both cases, we choose the (possibly trivial) prefix $\mathsf{sg}_{j}$ as a history starting in $\mathsf{first}(\rho^{(j)})$ that uses only vertices occurring in $\rho^{(j)}$ such that $\pi^{(j)}=\mathsf{sg}_{j}\cdot\mathsf{sg}_{j}^{\prime}$ is a simple play or a simple lasso. All states that occur infinitely often in $\pi^{(j)}$ are taken from $\rho^{(j)}{r_{j}}$. Therefore, the definition of $r_{j}$ ensures that $\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j)})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi^{(j)})$. Furthermore, Theorem 4.1 ensures that $\pi^{(j)}$ is an NE outcome (all vertices occurring in $\pi^{(j)}$ occur in the NE outcome $\rho^{(j)}$). This shows that $\pi^{(j)}$ satisfies all required properties of the induction that are unrelated to ii.
If $\pi^{(j)}$ satisfies ii, then we end the induction. Otherwise, we let $\rho^{(j+1)}$ witnessing that ii does not hold for $\pi^{(j)}$. To end the proof, it remains to show that $\rho^{(j+1)}$ is an NE outcome as the strict inclusion $\mathsf{SatPl}_{\mathcal{G}}(\pi^{j})\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j+1)})$ follows from the choice of $\rho^{(j+1)}$. This follows from Theorem 4.1, as all vertices occurring in $\rho^{(j+1)}$ occur in $\pi^{j}$ and $\mathsf{SatPl}_{\mathcal{G}}(\pi^{j})\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j+1)})$.

### v3

@@ 7.2.2 Outcomes without an infinitely occurring vertex
We now study NE outcomes in which no vertices occur infinitely often. These outcomes are specific to infinite arenas. We first present how to NE outcomes with a decomposition compatible with our approach from general NE outcomes. We obtain NE outcomes with an aperiodic infinite decomposition. We then present a variant of the previously used two-phase mechanism for these outcomes. While we leave the first phase is unchanged, finite memory no longer suffices to implement a decomposition-based approach for the second phase. Instead, we use a looser (but sufficient) variant of decomposition-based strategies to construct NEs.
@@
We first provide a counterpart to Lemma 7.5 for this case. Let $\pi^{\prime}\in\mathsf{Plays}(\mathcal{A})$ be the outcome of an NE from $v_{0}\in V$ such that no vertex occurs infinitely often in $\pi^{\prime}$. We observe that $\pi^{\prime}$ need not be a simple play: the objectives satisfied by $\pi$ may require visiting targets that are not connected by a simple play or a lasso. Therefore, simple plays do not suffice for our purpose, and we need memory to construct the outcome (even without taking profitable deviations into account).
We construct NE outcomes that can be obtained with two memory states. More precisely, we derive an NE outcome $\pi$ from $\pi^{\prime}$ that has an infinite simple segment decomposition $(\mathsf{sg}_{0},\mathsf{sg}_{1},\ldots)$ such that $\mathsf{SatPl}_{\mathcal{G}}(\pi)=\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})$, for all $j\neq j^{\prime}$ with the same parity, no vertex of $\mathsf{sg}_{j}$ occurs in $\mathsf{sg}_{j^{\prime}}$ and, for all $j\geq 1$, no targets of players whose objective is not satisfied by $\pi^{\prime}$ occurs in $\mathsf{sg}_{j}$. The condition on segments with the same parity implies that $\pi$ is the outcome of a strategy profile with memory $\{1,2\}$: intuitively, memory state $1$ is used when following an odd segment and $2$ is used when following an even segment.
@@
We define a finite-memory NE from $v_{0}$ whose outcome is $\pi$ via the two-phase mechanism previously used in the proof of Theorem 7.7. The first phase is defined identically to the previous section (see the proof of Theorem 7.7). For the second phase, we use a variation of the Mealy machines described in Definition 5.13. In said definition, for each segment index $j$ and each memory state of the form $(\mathcal{P}_{i},j)$ , we refer to the simple segment $\mathsf{sg}_{j}$ to define memory updates and next moves. This approach no longer yields a finite-memory strategy with the infinite decomposition $\mathcal{D}$. Instead of having a group of memory states for each segment, we use a group of memory states for each parity. More precisely, we use memory states of the form $(\mathcal{P}_{i},1)$ for all odd segments and memory states of the form $(\mathcal{P}_{i},2)$ for all even segments besides $\mathsf{sg}_{0}$ – the second component of the memory state indicates the parity of the current segment. When the end of an odd (resp. even) segment is reached in a memory state of the form $(\mathcal{P}_{i},1)$ (resp. $(\mathcal{P}_{i},2)$), the memory is updated to a state of the form $(\mathcal{P}_{i^{\prime}},2)$ (resp. $(\mathcal{P}_{i},1)$). We obtain a well-defined next-move function because segments of $\mathcal{D}$ with the same parity traverse pairwise disjoint sets of vertices.
@@
The update function $\mathsf{up}$ is defined as follows.
  1. (a)
    For all $\ell<r$, we let $\mathsf{up}(v_{\ell},v_{\ell})=v_{\ell+1}$.
  2. (b)
    For all $\ell\leq r$ and $v\neq v_{\ell}$,
    • if $\ell\geq 1$ and there exists $i\in I$ such that $v_{\ell-1}\in V_{i}$ (i.e., $v\in V_{I}$), we let $\mathsf{up}(v_{\ell},v)=\mathcal{P}_{i}$;
    • otherwise, $\mathsf{up}(v_{\ell},v)$ is left arbitrary.
  3. (c)
    For all states of the form $(\mathcal{P}_{i},p)\in M$ and all $v\in S_{p}$, we let $\mathsf{up}((\mathcal{P}_{i},p),v)=(\mathcal{P}_{i^{\prime}},p^{\prime})$ where
    • $\mathcal{P}_{i^{\prime}}$ is the player controlling $v$ if $v\in V_{I}$ and otherwise, $i^{\prime}=i$, and
    • $p^{\prime}=p$ if $v\notin L_{p}$, and otherwise we set $p^{\prime}=3-p$ (i.e., if $p=1$, it becomes $2$ and vice-versa).
  4. (d)
    For all states of the form $(\mathcal{P}_{i},p)\in M$ and all $v\in V\setminus S_{p}$, we let $\mathsf{up}((\mathcal{P}_{i},p),v)=\mathcal{P}_{i}$.
  5. (e)
    We let $\mathsf{up}(v_{r},v_{r})=\mathsf{up}((\mathcal{P}_{i},1),v_{r})$ for any $i\in I$ (the choice of $i$ here does not impact the following argument because $v_{r}\in S_{1}$).
  6. (f)
    Finally, for all $i\in I$ and $v\in V$, we let $\mathsf{up}(\mathcal{P}_{i},v)=\mathcal{P}_{i}$.
Let $i\in I$. We now define $\mathsf{nxt}_{i}$. Let $v\in V_{i}$.
  1. (a)
    We first consider memory states of the form $v_{\ell}$. Fix $\ell\leq r$.
    • If $v=v_{\ell}$ and $\ell\neq r$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=v_{\ell+1}$.
    • If $v=v_{\ell}$ and $\ell=r$, we let $\mathsf{nxt}_{i}(v_{r},v_{r})$ be the second vertex of $\mathsf{sg}_{1}$ (which exists because $\mathsf{sg}_{1}$ is non-trivial).
    • We distinguish two cases whenever $v\neq v_{\ell}$:
      • if $\ell\geq 1$ and there exists $i^{\prime}\in I$ such that $v_{\ell-1}\in V_{i^{\prime}}$ and $i^{\prime}\neq i$, we let $\mathsf{nxt}_{i}(v_{\ell},v)=\tau_{-i^{\prime}}(v)$;
      • we let $\mathsf{nxt}_{i}(v_{\ell},v)$ be arbitrary in all other cases.
  2. (b)
    We now deal with memory states of the form $(\mathcal{P}_{i^{\prime}},p)$. Fix $i^{\prime}\in I$ and $p\in\llbracket{}2\rrbracket{}$.
@@
    • If $v\in S_{p}\setminus L_{p}$, we set $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ to the vertex following $v$ in $\mathsf{sg}_{j}$, where $j\in 2\mathbb{N}+p$ is such that $v$ occurs in $\mathsf{sg}_{j}$ (such a $j$ is unique because all segments with the same parity traverse disjoint sets of vertices).
    • If $v\in L_{p}$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ be the second vertex of $\mathsf{sg}_{j}$ where $j\in 2\mathbb{N}+(3-p)$ is such that $v=\mathsf{first}(\mathsf{sg}_{j})$ ( such a $j$ is unique and $\mathsf{sg}_{j}$ has at least two vertices because it is non-trivial).
    • If $v\notin S_{p}$ and $i^{\prime}\neq i$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)=\tau_{-i^{\prime}}(v)$.
    • If $v\notin S_{p}$ and $i^{\prime}=i$, we let $\mathsf{nxt}_{i}((\mathcal{P}_{i^{\prime}},p),v)$ be arbitrary.
  3. (c)
    Finally, we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)=\tau_{-i^{\prime}}(v)$ if $i^{\prime}\neq i$, and otherwise we let $\mathsf{nxt}_{i}(\mathcal{P}_{i^{\prime}},v)$ be arbitrary.
@@ 7.3.1 Well-structured Nash equilibrium outcomes
In co-Büchi games, these deviations can be harmful: if we use a fully decomposition-based approach to define strategies from NE outcomes, plays resulting from profitable in-segment deviations may violate some objectives satisfied by the original NE outcome (Example 7.1). This motivates the use of a two-phase approach in co-Büchi games: we adopt a decomposition-based behaviour once in-segment deviations can no longer be harmful to winning players. Formally, we show that from any NE outcome $\pi^{\prime}$ in $\mathcal{G}$, we can derive an NE outcome $\pi$ from $\mathsf{first}(\pi^{\prime})$ that is a simple play or a simple lasso such that $\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi)$, and there are no plays $\pi^{\prime\prime}$ from $\mathsf{first}(\pi{r^{\star}})$ such that $\mathsf{SatPl}_{\mathcal{G}}(\pi)\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\pi^{\prime\prime})$ and $\pi^{\prime\prime}$ uses only vertices appearing in $\pi{r^{\star}}$ where $r^{\star}$ is the smallest $r\in\mathbb{N}$ such that no targets of players in $\mathsf{SatPl}_{\mathcal{G}}(\pi)$ appear in $\pi{r}$. The prefix $\pi{r^{\star}}$ corresponds to the first phase and the suffix $\pi{r^{\star}}$ to the second phase of our approach, and the condition imposed on $\pi$ guarantees the absence of profitable deviations. To construct $\pi$, we proceed by induction: we repeatedly simplify $\pi^{\prime}$ into well-chosen simple lasso or simple plays until we obtain a suitable play. The proof of the following lemma provides a formal construction.
@@
Assume that $\rho^{(j)}=v_{0}v_{1}\ldots$ is well-defined and let $r_{j}=\min\{r\in\mathbb{N}\mid\forall\ell\geq r,\,\forall i\in\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j)}),\,v_{\ell}\notin T_{i}\}$. We define $\pi^{(j)}$ as a combination of a history $\mathsf{sg}_{j}$ and a play $\mathsf{sg}_{j}^{\prime}$. We distinguish two cases for the suffixes. If $\rho^{(j)}{r_{j}}$ is a simple play, we let $\mathsf{sg}_{j}^{\prime}=\rho^{(j)}{r_{j}}$. Otherwise, we let $\mathsf{sg}_{j}^{\prime}$ be a simple lasso from $\mathsf{first}(\rho^{(j)}{r_{j}})$ such that all states that occur in $\mathsf{sg}_{j}^{\prime}$ occur in $\rho^{(j)}{r_{j}}$ (such a lasso exists because $\rho^{(j)}{r_{j}}$ is not a simple play). For both cases, we choose the (possibly trivial) prefix $\mathsf{sg}_{j}$ as a history starting in $\mathsf{first}(\rho^{(j)})$ that uses only vertices occurring in $\rho^{(j)}$ such that $\pi^{(j)}=\mathsf{sg}_{j}\cdot\mathsf{sg}_{j}^{\prime}$ is a simple play or a simple lasso. All states that occur infinitely often in $\pi^{(j)}$ are taken from $\rho^{(j)}{r_{j}}$. Therefore, the definition of $r_{j}$ ensures that $\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j)})\subseteq\mathsf{SatPl}_{\mathcal{G}}(\pi^{(j)})$. Furthermore, Theorem 4.1 ensures that $\pi^{(j)}$ is an NE outcome (all vertices occurring in $\pi^{(j)}$ occur in the NE outcome $\rho^{(j)}$). This shows that $\pi^{(j)}$ satisfies all required properties of the induction that are unrelated to ii.
If $\pi^{(j)}$ satisfies $i$, then we end the induction. Otherwise, we let $\rho^{(j+1)}$ witnessing that ii does not hold for $\pi^{(j)}$. To end the proof, it remains to show that $\rho^{(j+1)}$ is an NE outcome as the strict inclusion $\mathsf{SatPl}_{\mathcal{G}}(\pi^{j})\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j+1)})$ follows from the choice of $\rho^{(j+1)}$. This follows from Theorem 4.1, as all vertices occurring in $\rho^{(j+1)}$ occur in $\pi^{j}$ and $\mathsf{SatPl}_{\mathcal{G}}(\pi^{j})\subsetneq\mathsf{SatPl}_{\mathcal{G}}(\rho^{(j+1)})$.